The length is five units more than the width. The permeter is 9 times thw width. what is length and width?
Answer by robertb(5830)

l = w+5,
Perimeter is 2l+2w, and so 2l+2w = 9w. By direct substitution,
2(w+5)+2w = 9w.
2w+10+2w = 9w,
4w+10 = 9w,
10 = 5w,
w = 2, the width, while the length is 7.